Circular No. 8124

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003dv IN UGC 9638 R. Kotak and W. P. S. Meikle, Imperial College, London, on behalf of

the European Supernova Collaboration, report that spectra (resolution 0.27 nm) of the possible supernova reported on IAUC 8121, obtained at the Isaac Newton Telescope (+ IDS spectrograph) by P. Rodriguez-Gil on Apr. 26.11 (range ∼ 500–750 nm) and Apr. 27.06 UT (range ∼ 350–570 nm), indicate that it is of type IIn, ∼ 1 week past explosion. The spectrum comprises a blue continuum upon which a number of emission features are superposed. The continuum appears to show a break at ∼ 530 nm, blueward of which it becomes steeper. The spectrum lacks broad P-Cyg-type profiles and bears a striking resemblance to early-epoch spectra of the type-IIn SN 1998S (IAUC 6829; Leonard et al. 1999, Ap.J. 536, 239; Fassia et al. 2001, MNRAS 325, 907). Balmer lines in emission are present and exhibit a multicomponent structure. The Hα profile comprises a narrow but resolved component (FWHM ∼ 350 km/s) superposed on a much broader profile of FWZI ∼ 10000 km/s. The narrow component appears to lie slightly redward (∼ 350 km/s) of the peak of the broad component. A broad emission feature at ∼ 470 nm is present, which is attributed to a blend of He II 468.6-nm and C III 464.8-nm/N III 464.0-nm (cf. SN 1998S; IAUC 6832). A weak He I 587.6-nm emission feature may also be present. In SN 1998S, the early-time emission features yielded valuable information about the interaction of the supernova with the ‘circumstellar medium’ of the progenitor star, but these features persisted for barely 10 days. High-resolution spectral coverage over the next few days is urged.

SUPERNOVA 2003dw IN MCG +10-24-51 D. Rich, Hampden, ME, reports his discovery of an apparent supernova

(mag ∼ 16.7) on unfiltered CCD frames taken with a 0.31-m reflector on Apr. 26.17 and 28.10 UT. SN 2003dw is located at α = 16^{h}45^{m}51^{s}.10, δ = +59^{o}37′38′′.0 (equinox 2000.0), which is ∼ 11′′.4 west and 19′′.0 north of the center of MCG +10-24-51. The new object is not visible on the red/blue Palomar Sky Survey digital images or on CCD frames taken by Rich on 2002 Apr. 5.32 and May 22.24 (limiting mag 18.0).

V4745 SAGITTARII = NOVA SAGITTARII 2003 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4745 Sgr has been given to this object (cf. IAUC 8123). Corrigendum. On IAUC 8123, line 2, for Nicolas read Nicholas

2003 April 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green